How much 10.0 M HNO3 must be added to 1.00 L of a buffer that is 0.0100 M acetic acid and 0.100 M sodium acetate to reduce the pH to 4.80 ?
1
Expert's answer
2019-10-29T09:11:07-0400

pH=pKa+lg(([NaAc]-x)/([HAc]+x))

4.8=4.76+lg((0.1-x)/(0.01+x))

lg((0.1-x)/(0.01+x))=0.04

(0.1-x)/(0.01+x)=1.096478

x=0.0425

V=1000*x/C=1000*0.425/10=4.25 (ml)

Answer: 4.25 ml.